OpenStudy (anonymous):

An acute angle θ is in a right triangle with sin θ =6/7 . What is the value of cot θ

OpenStudy (anonymous):



OpenStudy (anonymous):

How do I find the adjacent side?

OpenStudy (phi):

draw a triangle , label theta and the legs that have length 6 and 7 figure out the 3rd side (use pythagoras a^2+b^2= c^2) what is the defintion of cot ? 1/tan and tan is opp/adj

OpenStudy (phi):

I assume you can find the 3rd leg: 6^2 + b^2 = 7^2 b^2= 7^2 - 6^2 b^2= 49-36 b^2= 13 b= sqrt(13)

OpenStudy (anonymous):

I got it thanks